# Magneto-optical transport in tilted type-I double-Weyl semimetals in the presence of orbital magnetic moment

### v1

@@ III.3 Calculatons of quadractic-B contributon to the conductivity $\sigma_{ab}^{(B^{2})}$
It is noted that all the other magnetoconductivity components are zero. It is clear that the above conductivity equations are independent of chirality, i.e., the Weyl cones with opposite chiralities have the same contribution to the conductivity. In contrast to single WSMs, only the conductivity component $\sigma_{zz}^{(B^{2},m)}(\omega)$ is always negative, a result that will leave the total conductivity weakened. Other components of conducivities are enhanced due to orbital magnetic moment.
@@ III.4 Hall conductivities $\sigma_{ab}^{(H,0)}$ and $\sigma_{ab}^{(H,B)}$
The intrinsic Hall effect is engendered by the Berry curvature, as presented by the second term of Eq. (III), for which the further calculation gives
@@
$\sigma_{ab}^{(H,0)}(\omega)$ $=$ $\frac{e^{2}}{\hbar(2\pi)^{3}}\epsilon_{abc}\int d^{3}k\Omega_{c}^{s}(\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B})\Bigl{(}-\frac{\partial f_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\Bigr{)}$ (48)
where $\epsilon_{abc}$ is the Levi-Civita symbol with $a,b,c,\in{x,y,z}$ The first term $\sigma_{ab}^{(H,0)}(\omega)$ in Eq. (46), referring to the anomalous Hall effect, is not equivalent to zero only in the system with broken time-reversal symmetry [22]. While the second term $\sigma_{ab}^{(H,0)}(\omega)$stands for the ordinary Hall conductivity linear in B, which is counterpart to a semiclassical description related to Landau level formation in the quantum limit.
@@
where $\sigma_{1}^{H}=-\frac{e^{2}}{\hbar}\frac{Be}{4\pi}\frac{(-2+t_{s}^{2}+\sqrt{1-t_{s}^{2}})}{t_{s}^{4}\mathit{v}_{F}\hbar}\alpha_{2}$
In contrast to the single WSM node, the B-linear contribution to the Hall conductivity of Weyl semimetals depends on the tilt and depends on the chirality of the valley as in the case of single WSM. The linear Hall conductivity is independent of $\mu$ while in case of single WSM, it is inversely proportional to $\mu$.
IV Second order non-linear response of DWSMs
Now, we exlore the second-order nonlinear magneto-optical response of Weyl semimetals. Substituting Eq.(9) into Eq.(10), neglecting the Lorentz force term, and retaining terms up to second order in E, we obtain
@@ IV Second order non-linear response of DWSMs
$\tilde{f}_{2}^{s}=\frac{e\tau^{2}}{(1-i\omega\tau)(1-2i\omega\tau)}\biggl{\{}\textbf{E}\cdot\frac{\partial}{\partial\bm{k}}\biggr{[}\Bigl{(}e\textbf{E}+\frac{e^{2}}{\hbar}(\textbf{E}\cdot\textbf{B})\Omega_{\bm{k}}^{s}-\frac{e^{2}}{\hbar}(\textbf{B}\cdot\Omega_{\bm{k}}^{s})\textbf{E}\Bigr{)}\cdot\bm{\mathit{v}}_{\bm{k}}^{s}\frac{\partial f_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}-\frac{e}{\hbar}\textbf{E}\cdot\frac{\partial}{\partial\bm{k}}\biggl{(}\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B}\frac{\partial f_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\biggr{)}\biggr{]}$
$+\biggl{[}\frac{e^{2}}{\hbar}(\textbf{E}\cdot\textbf{B})\Omega_{\bm{k}}^{s}-\frac{e^{2}}{\hbar}(\textbf{B}\cdot\Omega_{\bm{k}}^{s})\textbf{E}\biggr{]}\cdot\frac{\partial}{\partial\bm{k}}\biggl{(}\textbf{E}\cdot\bm{\mathit{v}}_{\bm{k}}^{s}\frac{\partial f_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\biggr{)}\biggr{\}}$ (52)
@@
$\bm{j}_{1}=-\frac{e}{(2\pi)^{3}}\int d^{3}k\Bigl{[}\bm{\mathit{\tilde{v}}}_{\bm{k}}^{s}+\frac{e}{\hbar}(\Omega_{\bm{k}}^{s}\cdot\bm{\mathit{\tilde{v}}}_{\bm{k}}^{s})\textbf{B}\Bigr{]}\tilde{f}_{2}^{s}$
$-\frac{e^{2}}{2\pi)^{3}\hbar}\int d^{3}k\textbf{E}\times\Omega_{\bm{k}}^{s}\tilde{f}_{1}^{s}$ (53)
@@
In this subsection, we calculate the second harmonic current of the Weyl semimetals system in absence of magnetic fields $\textbf{B}=0$. Inserting Eq. (IV) into the first term of Eq. (IV), the second harmonic conductivity tensor can be written as
@@
Except the aaz, aza, and zaa (a = x, y, z) components of the second harmonic conductivity tensor are nonzero and all other components equal to zero [22, 30]
$\sigma_{zzz}^{0}(2\omega)=-\frac{3\pi\mathit{v}_{F}^{2}}{8\alpha_{2}}\sigma_{DL}\Bigl{[}\frac{2-2\sqrt{1-t_{s}^{2}}}{t_{s}^{3}}+\frac{-3+2\sqrt{1-t_{s}^{2}}}{t_{s}}\Bigr{]}$ (56)
$\sigma_{xxz}^{0}(2\omega)=\sigma_{DL}\Bigl{[}\frac{3-2t_{s}^{2}}{t_{s}^{3}}+3\frac{1-t_{s}^{2}}{2t_{s}^{4}}\ln\frac{1-t_{s}}{1+t_{s}}\Bigr{]}$ (57)
where $\sigma_{DL}=\frac{e^{3}\tau^{2}}{\hbar(1-2\omega t)(1-i\omega t)6\pi^{2}\hbar}$ is the Drude-like frequency dependent complex conductivity.
The second harmonic conductivity tensor satisfies the relation $\sigma_{xxz}^{0}(2\omega)\sigma_{yyz}^{0}(2\omega)=\sigma_{zxx}^{0}(2\omega)=\sigma_{zyy}^{0}(2\omega)=\sigma_{xzx}^{0}(2\omega)=\sigma_{yzy}^{0}(2\omega)$,$\sigma_{xxx}^{0}(2\omega)=\sigma_{yyy}^{0}(2\omega)=0$ and it does not depend on the chirality of Weyl node. The total second harmonic conductivity in tilted Weyl semimetals is the sum of a pair of Weyl nodes. For the case with tilt inversion symmetry $t_{+}=-t_{-},\sigma_{abc}(2\omega)=0$. For the case with broken tilt inversion symmetry $t_{+}=t_{-},\sigma_{abc}(2\omega)\neq 0$.
@@
Inserting Eq. (IV) into the first term of Eq. (IV), we write the lnear B dependent second harmonic conductivity tensor
$\sigma_{abc}^{B}(2\omega)=\frac{\tau^{2}}{(1-i\omega\tau)(1-2i\omega\tau)}\frac{e^{3}}{\hbar(2\pi)^{3}}\int d^{3}k\biggl{\{}\frac{\partial\mathit{v}_{a}^{s}}{\partial k_{c}}\Bigl{[}-\frac{eB_{b}}{\hbar}(\Omega_{\bm{k}}^{s}\cdot\bm{\mathit{v}}_{\bm{k}}^{s})+\frac{e\mathit{v}_{b}^{s}}{\hbar}(\Omega_{\bm{k}}^{s}\cdot\bm{B})\Bigr{]}-\frac{eB_{c}}{\hbar}\frac{\partial}{\partial\bm{k}}\cdot(\mathit{v}_{a}^{s}\Omega_{\bm{k}}^{s})\mathit{v}_{b}^{s}$
$+\frac{e}{\hbar}\frac{\partial[\mathit{v}_{a}^{s}(\Omega_{\bm{k}}^{s}\cdot\bm{B})]}{\partial k_{c}}\mathit{v}_{b}^{s}-\frac{eB_{a}}{\hbar}\frac{\partial(\Omega_{\bm{k}}^{s}\cdot\bm{\mathit{v}}_{\bm{k}}^{s})}{\partial k_{c}}\mathit{v}_{b}^{s}+\frac{\partial^{2}(\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B})}{\hbar\partial k_{a}\partial k_{c}}\mathit{v}_{b}^{s}-\frac{\partial^{2}\mathit{v}_{a}^{s}}{\hbar\partial k_{a}\partial k_{c}}(\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B})\biggr{\}}\Bigl{(}-\frac{\partial f_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\Bigr{)}$ (58)
We will explore Eq.(IV.2) with follwing two cases:
@@
$\sigma_{xzx}^{B}(2\omega)=2s\sigma_{2}^{(B)}\frac{\sqrt{1-t_{s}^{2}}}{t_{s}^{4}}[2-t_{s}^{2}-2\sqrt{1-t_{s}^{2}}]$ (59)
$\sigma_{zxx}^{B}(2\omega)=-s\sigma_{2}^{(B)}\frac{1}{t_{s}^{4}}[2-t_{s}^{2}-2\sqrt{1-t_{s}^{2}}]$ (60)
where $\sigma_{2}^{(B)}=\frac{e^{3}\tau^{2}\alpha_{2}}{\hbar(1-2\omega t)(1-i\omega t)}\frac{eB}{4\pi\hbar}$
We can check that other nonzero second harmonic conductivity components satisfy the relations: $\sigma_{zxx}^{B}(2\omega)=\sigma_{zyy}^{B}(2\omega)=\sigma_{xzx}^{B}(2\omega)=\sigma_{yzy}^{B}(2\omega)$. In contrast to single WSM, the B-linear contribution to the second harmonic conductivity deends on the tilt. Summing the conductivity over the Weyl cones with opposite chirality cancels each other, leading to the disappearance of the total B-linear contribution to the second harmonic conductivity.

### v2

@@ III.3 Calculations of quadratic-B contribution to the conductivity $\sigma_{ab}^{(B^{2})}$
It is noted that all the other magnetoconductivity components are zero. It is clear that the above conductivity equations are independent of chirality, i.e., the Weyl cones with opposite chiralities have the same contribution to the conductivity. The conductivity component $\sigma_{zz}^{(B^{2},m)}(\omega)$ is always negative and follow a power laws condition $\sigma_{zz}^{(B^{2},m)}(\omega)\propto J$, a result that will suppress the total conductivity compare to its Berry curvature parts.[see Fig(7)] .
In the present system, the planar Hall effect can take place [34, 35, 36, 33, 37, 38] and manifest itself in a nonzero conductivity $\sigma_{xy}^{(B^{2})}(\omega)$. Similar to the diagonal component of the conductivity, $\sigma_{xy}^{(B^{2})}(\omega)$ also consists of the contributions from the Berry curvature and the orbital magnetic moment. Figure (8) shows an effect of the tilt on the planar Hall magnetoconductivity. It is seen that the total planar Hall conductivity is suppressed in the case of single WSM when the orbit magnetic moment is present as shown in Fig.(8a). On the other hand, the total conductivity is enhanced with tilt parameter due to positive magnetic orbital moment for double and triple WSMs as shown in Fig.(8b) and Fig.(8c). Figure (9) shows the planar Hall magnetoconductivity as a function of the THz incident light at $\gamma=\pi/4$. The real and imaginary parts of total conductivities are suppressed in the case of single WSM due to negative contribution of real and imagnary parts of magnetic orbital moment conductivtes. For double and triple WSMs, the real and imagnary parts of conductivities are enhanced due to positive contributions of real and imaginary parts of orbital magnetic moment conductivites.
The $B^{2}$ dependence of the magnetoconductivity has been observed experimentally in the materials such as GdPtBi and TaP [37, 38]. TaS and GdPtBi have isotropic and quadratic band dispersions respectively. Therefore they are examples of single and double WSMs. Recently, it is also shown that in the materials GdPtBi, a very strong planar Hall effect has been reported, which is due to the Berry curvature and chiral anomaly contributions [37]. Besides chiral anomaly, the planar Hall effect may be induced by the orbital magnetic moment.
@@ III.4 Hall conductivities $\sigma_{ab}^{(H,0)}$ and $\sigma_{ab}^{(H,B)}$
The intrinsic Hall effect is engendered by the Berry curvature, as presented by the second term of Eq. (17), for which further calculation gives
@@
$\sigma_{ab}^{(H,0)}(\omega)$ $=$ $\frac{e^{2}}{\hbar(2\pi)^{3}}\epsilon_{abc}\int d^{3}k\Omega_{c}^{s}(\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B})\Bigl{(}\frac{\partial f_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\Bigr{)}$ (86)
where $\epsilon_{abc}$ is the Levi-Civita symbol with $a,b,c,\in{x,y,z}$ The first term $\sigma_{ab}^{(H,0)}(\omega)$ in Eq. (84), referring to the anomalous Hall effect, is not equivalent to zero only in the system with broken time-reversal symmetry [27, 22]. While the second term $\sigma_{ab}^{(H,0)}(\omega)$stands for the ordinary Hall conductivity linear in B, which is the counterpart to a semiclassical description related to Landau level formation in the quantum limit [32].
@@
where
for J=1
$\sigma_{1}^{H}=-\frac{e^{3}}{\hbar}\frac{B\mathit{v}_{F}}{24\pi^{2}\mu}$ (88)
for J=2
$\sigma_{1}^{H}=\frac{e^{2}}{\hbar}\frac{Be}{4\pi}\frac{(-2+t_{s}^{2}+2\sqrt{1-t_{s}^{2}})}{t_{s}^{4}\mathit{v}_{F}\hbar}\alpha_{2}$ (89)
for J=3
$\begin{split}\sigma_{1}^{H}=&-\frac{{e^{3}}B\alpha_{3}^{\frac{2}{3}}\mu^{\frac{1}{3}}}{\hbar^{\frac{7}{3}}{4\pi^{\frac{3}{2}}}\mathit{v}_{F}}\frac{(3)^{\frac{1}{2}}\Gamma{(\frac{5}{6})}}{(1-t_{s})^{\frac{4}{3}}t_{s}^{3}\Gamma{(\frac{4}{3})}}\Bigg{[}(-1+t_{s})(-35+t_{s}(5+3t_{s})){}_{2}F_{1}{\bigg{(}\frac{-2}{3},\frac{5}{3},\frac{7}{3},\frac{2t_{s}}{-1+t_{s}}\bigg{)}}+\\ &\big{(}-35+t_{s}(-10+t_{s}(8+t_{s}))\big{)}{}_{2}F_{1}{\bigg{(}\frac{1}{3},\frac{5}{3},\frac{7}{3},\frac{2t_{s}}{-1+t_{s}}\bigg{)}}\Bigg{]}\end{split}$ (90)
So, we can rewrite the above components follow the power laws $\sigma_{1}^{H}\propto\frac{\alpha_{J}^{\frac{2}{J}}\mu^{(1-\frac{2}{J})}}{\hbar^{2(1-\frac{1}{J})}}$.
In contrast to the single WSM node, the B-linear contribution to the Hall conductivity of multi-WSMs depends on the tilt parameter. The linear Hall conductivity is independent of $t_{s}$ in the case of a single WSM while in the case double and triple cases, its depend on $t_{s}$ as shown in Fig.(10).
Refer to caption  Refer to caption  Refer to caption
Figure 10: The dependence of the optical conductivity on the tilt $t_{+}=0.5$ for the case of B $\parallel t_{s}$. for (a)single WSM (b) double WSM and (c) triple WSM. The other parameters are the same as those of Fig.(1)
IV Second order non-linear response of multi-WSMs
Now, we explore the second-order nonlinear magneto-optical response of Weyl semimetals. Substituting Eq.(9) into Eq.(10), neglecting the Lorentz force term, and retaining terms up to second order in E, we obtain
@@ IV Second order non-linear response of multi-WSMs
$\tilde{f}_{2}^{s}$ $=$ $\frac{e\tau^{2}}{\hbar(1-i\omega\tau)(1-2i\omega\tau)}\biggl{\{}\textbf{E}\cdot\frac{\partial}{\partial\bm{k}}\biggr{[}\Bigl{(}e\textbf{E}+\frac{e^{2}}{\hbar}(\textbf{E}\cdot\textbf{B})\Omega_{\bm{k}}^{s}-\frac{e^{2}}{\hbar}(\textbf{B}\cdot\Omega_{\bm{k}}^{s})\textbf{E}\Bigr{)}\cdot\bm{\mathit{v}}_{\bm{k}}^{s}\frac{\partial f_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}-\frac{e}{\hbar}\textbf{E}\cdot\frac{\partial}{\partial\bm{k}}\biggl{(}\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B}\frac{\partial f_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\biggr{)}\biggr{]}$ (93)
$+$ $\biggl{[}\frac{e^{2}}{\hbar}(\textbf{E}\cdot\textbf{B})\Omega_{\bm{k}}^{s}-\frac{e^{2}}{\hbar}(\textbf{B}\cdot\Omega_{\bm{k}}^{s})\textbf{E}\biggr{]}\cdot\frac{\partial}{\partial\bm{k}}\biggl{(}\textbf{E}\cdot\bm{\mathit{v}}_{\bm{k}}^{s}\frac{\partial f_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\biggr{)}\biggr{\}}$
@@
$\bm{j}_{1}$ $=$ $-\frac{e}{(2\pi)^{3}}\int d^{3}k\Bigl{[}\bm{\mathit{\tilde{v}}}_{\bm{k}}^{s}+\frac{e}{\hbar}(\Omega_{\bm{k}}^{s}\cdot\bm{\mathit{\tilde{v}}}_{\bm{k}}^{s})\textbf{B}\Bigr{]}\tilde{f}_{2}^{s}$ (94)
- $\frac{e^{2}}{2\pi)^{3}\hbar}\int d^{3}k\textbf{E}\times\Omega_{\bm{k}}^{s}\tilde{f}_{1}^{s}$
@@ IV.1 Second harmonic conductivity $\sigma_{abc}^{0}$ in the absence of magnetic field
In this subsection, we calculate the second harmonic current of the Weyl semimetals system in the absence of magnetic fields $\textbf{B}=0$. Inserting Eq. (93) into the first term of Eq. (94), the second harmonic conductivity tensor can be written as
@@
Except for the aaz, aza, and zaa (a = x, y, z) components of the second harmonic conductivity tensor are nonzero and all other components equal to zero [22, 39]
for J=1
$\sigma_{zzz}^{0}(2\omega)$ $=$ $\frac{\mu}{\pi\hbar^{2}}\sigma_{DL}\Bigl{[}\frac{-6+4t_{s}^{2}}{t_{s}^{3}}-\frac{3(1-t_{s}^{2})}{t_{s}^{4}}\ln\frac{1-t_{s}}{1+t_{s}}\Bigr{]}$
$\sigma_{xxz}^{0}(2\omega)$ $=$ $\frac{\mu}{2\pi\hbar^{2}}\sigma_{DL}\Bigl{[}\frac{6}{t_{s}^{3}}+\frac{3-t_{s}^{2}}{t_{s}^{4}}\ln\frac{1-t_{s}}{1+t_{s}}\Bigr{]}$ (98)
for J=2
$\sigma_{zzz}^{0}(2\omega)$ $=$ $\frac{\mathit{v}_{F}^{2}}{4\hbar\alpha_{2}}\sigma_{DL}\Bigl{[}\frac{2-2\sqrt{1-t_{s}^{2}}}{t_{s}^{3}}+\frac{-3+2\sqrt{1-t_{s}^{2}}}{t_{s}}\Bigr{]}$
$\sigma_{xxz}^{0}(2\omega)$ $=$ $\frac{\mu}{\pi\hbar^{2}}\sigma_{DL}\Bigl{[}\frac{6}{t_{s}^{3}}+\frac{3-t_{s}^{2}}{t_{s}^{4}}\ln\frac{1-t_{s}}{1+t_{s}}\Bigr{]}$ (100)
for J=3
$\sigma_{zzz}^{0}(2\omega)$ $=$ $\frac{\mathit{v}_{F}^{2}\sigma_{DL}}{2\pi^{\frac{1}{2}}\hbar^{\frac{2}{3}}\alpha_{3}^{\frac{2}{3}}\mu^{\frac{1}{3}}}\frac{-2\Gamma\big{(}\frac{4}{3}\big{)}}{3(1-t_{s})^{\frac{2}{3}}\Gamma{(\frac{11}{6})}}\bigg{[}(-1+t_{s}){}_{2}F_{1}{\Big{(}\frac{2}{3},\frac{4}{3},\frac{8}{3},\frac{2t_{s}}{-1+t_{s}}\Big{)}}+{}_{2}F_{1}{\Big{(}\frac{2}{3},\frac{7}{3},\frac{11}{3},\frac{2t_{s}}{-1+t_{s}}\Big{)}}\bigg{]}$ (101)
$\sigma_{xxz}^{0}(2\omega)$ $=$ $\frac{\mu}{2\pi\hbar^{2}}\sigma_{DL}\Bigl{[}\frac{18}{t_{s}^{3}}-\frac{3(-3+t_{s}^{2})}{t_{s}^{4}}\ln\Bigl{[}\frac{1-t_{s}}{1+t_{s}}\Bigr{]}\Bigr{]}$ (102)
where $\sigma_{DL}=\frac{e^{3}\tau^{2}}{(1-2i\omega\tau)(1-i\omega\tau)4\pi\hbar}$ is the Drude-like frequency dependent complex conductivity.
So, we can rewrite the above components follow the power laws $\sigma_{zzz}^{(0)}(2\omega)\propto\frac{\mu^{-(1-2/J)}}{\alpha_{J}^{\frac{2}{J}}\hbar^{\frac{2}{J}}}$ and $\sigma_{xxz}^{(0)}(2\omega)\propto\frac{J\mu}{\hbar^{2}}$.
Refer to caption  Refer to caption  Refer to caption
Figure 11: The nonlinear conductivities for the process of second harmonic generation of the tilt $t_{+}=0.5$, for (a) single WSM (b)double WSM and (c) triple WSM. The other parameters are the same as those of Fig.(1)
The second harmonic conductivity tensor satisfies the relation $\sigma_{zxx}^{0}(2\omega)=\sigma_{zyy}^{0}(2\omega)=\sigma_{xzx}^{0}(2\omega)=\sigma_{yzy}^{0}(2\omega)=\sigma_{xxz}^{0}(2\omega)=\sigma_{yyz}^{0}(2\omega)$, and it does not depend on the chirality of Weyl node. The total second harmonic conductivity in tilted Weyl semimetals is the sum of a pair of Weyl nodes. For the case with tilt inversion symmetry $t_{+}=-t_{-},\sigma_{abc}(2\omega)=0$. For the case with broken tilt inversion symmetry $t_{+}=t_{-},\sigma_{abc}(2\omega)\neq 0$.
It is noted to see from Fig.(11) that the $\sigma_{abc}^{0}(2\omega)$ becomes exactly zero when $t_{s}\rightarrow 0$. The presence of the finite tilt is needed to get the second harmonic generation in mWSMs. It is clear from Fig.(11) that $\sigma_{zzz}^{0}(2\omega)$ is more sensitive to tilt that $\sigma_{xxz}^{0}(2\omega)$. Fig.(13) shows the frquency variation of multi-WSMs.
@@ IV.2 Second harmonic conductivity $\sigma_{abc}^{B}$ in the presence of linear magnetic field
Inserting Eq. (93) into the first term of Eq. (94), we write the linear B dependent second harmonic conductivity tensor
$\sigma_{abc}^{B}(2\omega)$ $=$ $\frac{\tau^{2}}{(1-i\omega\tau)(1-2i\omega\tau)}\frac{e^{3}}{\hbar(2\pi)^{3}}\int d^{3}k\biggl{\{}\frac{\partial\mathit{v}_{a}^{s}}{\partial k_{c}}\Bigl{[}-\frac{eB_{b}}{\hbar}(\Omega_{\bm{k}}^{s}\cdot\bm{\mathit{v}}_{\bm{k}}^{s})+\frac{e\mathit{v}_{b}^{s}}{\hbar}(\Omega_{\bm{k}}^{s}\cdot\bm{B})\Bigr{]}-\frac{eB_{c}}{\hbar}\frac{\partial}{\partial\bm{k}}\cdot(\mathit{v}_{a}^{s}\Omega_{\bm{k}}^{s})\mathit{v}_{b}^{s}$ (103)
$+$ $\frac{e}{\hbar}\frac{\partial[\mathit{v}_{a}^{s}(\Omega_{\bm{k}}^{s}\cdot\bm{B})]}{\partial k_{c}}\mathit{v}_{b}^{s}-\frac{eB_{a}}{\hbar}\frac{\partial(\Omega_{\bm{k}}^{s}\cdot\bm{\mathit{v}}_{\bm{k}}^{s})}{\partial k_{c}}\mathit{v}_{b}^{s}+\frac{\partial^{2}(\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B})}{\hbar\partial k_{a}\partial k_{c}}\mathit{v}_{b}^{s}-\frac{\partial^{2}\mathit{v}_{a}^{s}}{\hbar\partial k_{a}\partial k_{c}}(\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B})\biggr{\}}\Bigl{(}-\frac{\partial f_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\Bigr{)}$
We will explore Eq.(103) with following two cases:
Refer to caption  Refer to caption  Refer to caption
Figure 12: The frequency dependence of optical conductivity at tilt $t_{+}=0.5$, for (a) single WSM (b)double WSM and (c) triple WSM. The other parameters are the same as those of Fig.(1)
@@
for J=1
$\sigma_{xzx}^{B}(2\omega)$ $=$ $\frac{\mathit{v}_{F}^{2}}{12\pi^{2}\mu}s\sigma_{2}^{(B)}$ (104)
$\sigma_{zxx}^{B}(2\omega)$ $=$ $-\frac{\mathit{v}_{F}^{2}}{12\pi^{2}\mu}s\sigma_{2}^{(B)}$ (105)
for J=2
$\sigma_{xzx}^{B}(2\omega)$ $=$ $\frac{\alpha_{2}}{2\pi}s\sigma_{2}^{(B)}\frac{\sqrt{1-t_{s}^{2}}}{t_{s}^{4}}[2-t_{s}^{2}-2\sqrt{1-t_{s}^{2}}]$
$\sigma_{zxx}^{B}(2\omega)$ $=$ $-\frac{\alpha_{2}}{2\pi}s\sigma_{2}^{(B)}\frac{1}{t_{s}^{4}}[2-t_{s}^{2}-2\sqrt{1-t_{s}^{2}}]$ (107)
for J= 3
$\sigma_{xzx}^{B}(2\omega)=\frac{2s\sigma_{2}^{(B)}\alpha_{3}^{\frac{2}{3}}\mu^{\frac{1}{3}}}{\hbar^{\frac{4}{3}}\pi^{\frac{1}{2}}}\frac{1}{{3(3)^{\frac{1}{2}}(1-t_{s})^{\frac{4}{3}}t_{s}^{3}{\big{(}\frac{1+t_{s}}{1-t_{s}}\big{)}^{\frac{2}{3}}}\Gamma{(\frac{1}{6})}\Gamma{(\frac{1}{3})}}}\Bigg{[}(-1+t_{s})\bigg{\{}-70+3t_{s}\big{\{}-10+t_{s}\big{(}11\\ +18t_{s}(-1+\bigg{(}\frac{1+t_{s}}{1-t_{s}}\bigg{)}^{\frac{2}{3}}\bigg{(}-1+\frac{2}{1+t_{s}}\bigg{)}^{\frac{2}{3}})\big{)}\big{\}}\bigg{\}}+7(-10+9t_{s}^{2}){{}_{2}F_{1}{\bigg{(}\frac{2}{3},1,\frac{7}{3},\frac{2t_{s}}{-1+t_{s}}\bigg{)}}}\Bigg{]}$ (108)
$\begin{split}\sigma_{zxx}^{B}(2\omega)=&\frac{-6s\sigma_{2}^{(B)}\alpha_{3}^{\frac{2}{3}}\mu^{\frac{1}{3}}}{\hbar^{\frac{4}{3}}\pi^{\frac{1}{2}}}\frac{1}{{81(3)^{\frac{1}{2}}(1-t_{s})^{\frac{4}{3}}t_{s}^{5}(\frac{1+t_{s}}{1-t_{s}})\Gamma{(\frac{1}{6})}\Gamma{(\frac{4}{3})}}}\\ &\Bigg{[}{(-1+t_{s})\bigg{\{}2912+3t_{s}\big{\{}416+3t_{s}\big{(}-293+t_{s}(-89+31t_{s})\big{)}\big{\}}\bigg{\}}+\bigg{(}2912-3885t_{s}^{2}+1080t_{s}^{4}\bigg{)}{{}_{2}F_{1}{\bigg{(}\frac{2}{3},1,\frac{7}{3},\frac{2t_{s}}{-1+t_{s}}}}\bigg{)}}\Bigg{]}\end{split}$ (109)
where $\sigma_{2}^{(B)}=\frac{e^{4}\tau^{2}B}{\hbar^{2}(1-2i\omega\tau)(1-i\omega\tau)}$
So, we can rewrite the above components follow the power laws $\sigma_{xzx}^{(B)}(2\omega)\propto\frac{\alpha_{J}^{\frac{2}{J}}\mu^{(1-\frac{2}{J})}}{\hbar^{2(1-\frac{1}{J})}}$ and $\sigma_{zxx}^{(B)}(2\omega)\propto\frac{\alpha_{J}^{\frac{2}{J}}\mu^{(1-\frac{2}{J})}}{\hbar^{2(1-\frac{1}{J})}}$.
We can check that other nonzero second harmonic conductivity components satisfy the relations: $\sigma_{zxx}^{B}(2\omega)=\sigma_{zyy}^{B}(2\omega)=\sigma_{xzx}^{B}(2\omega)=\sigma_{yzy}^{B}(2\omega)$. In contrast to single WSM, the B-linear contribution to the second harmonic conductivity depends on the tilt. Summing the conductivity over the Weyl cones with opposite chirality cancels each other, leading to the disappearance of the total B-linear contribution to the second harmonic conductivity.